OpenStudy (anonymous):

i want the sum of series: √2+√2+√2+√2+.........infinity and there is a big underroot and then inh and in that i cant draw

OpenStudy (anonymous):

\[let y=\sqrt{2+\sqrt{2+\sqrt{2........}}}\] \[y ^{2}=2+\sqrt{2+\sqrt{2+......}}\]\[y ^{2}=2+y\]\[y ^{2}-y-2=0\]gives you y=2,-1. where -1 is a redundant root came due to the squaring so y=2.